anyone can help me with it...? biochemistry

2. A buffer contains 0.01 mole of lactic acid (pKa-3.86) and 005 mole of sodium lactate per liter. (a) Calculate the pH of the buffer. (b) calculate the change in plH when 5 ml of 0.5 M HCl is added to 1L of the buffer. (c) what pll change would you expect if you added the same quantity of HCI to IL of pure water (8 points)?
Answer #1

(A) pH = 3.86 + log 0.050/ 0.010=4.56

(B)
moles H+ added = 5 x 10^-3 L x 0.5 M=0.0025

moles lactic acid = 0.010 + 0.0025=0.0125
moles lactate = 0.050 - 0.0025 = 0.0474
pH = 3.86 + log 0.0474/ 0.0125 =4.44
change = 4.56 - 4.44 =0.12

(C)

pH water = 7.0
Total volume = 1005 mL = 1.005 L
[H+]= 0.0025/ 1.005=0.00249 M
pH = 2.6
change = 7.0 - 2.6 =4.4